OpenStudy (anonymous):

Simplify (squareroot) -16 with imaginary unit i Thanks

OpenStudy (ranga):

\[\Large \sqrt{-16} = \sqrt{16 \times -1} = \sqrt{16} \times \sqrt{-1} = 4i\]

OpenStudy (ranga):

The definition of i is:\[\Large i = \sqrt{-1}\]

OpenStudy (anonymous):

thank you!

OpenStudy (ranga):

you are welcome.